solve for the reactions in each journal bearing below
Solve for the reactions in each Journal Bearing below. AM 400 N 0.6 m CB 600ND 0.6 m 0.6 m 0.4 m
Answer #1

Given data in the form of free body diagram which means removing all supports & representing their respective reactions on thconsidering moment equiliborum equatron Along x-directron, ie Mx=0 Moment acurs due perpendicular forces so consides the forcHere RBX as Now along y direction Applicable forces are Goon, Rex Myco - 600 XO 6 -Rex X0-4 = 0 - Rcx = -900N ans perpendiculfrom Raxt Rext Rex = 0 Rox+ 1400 + (-900) = 0 - Rox=-500N ans Reaction are RAX=250 ON RB X= 140ON RREX = -900N RAZ= 333.33 N